how much water must be evaporated from 64oz of 2% solution to make a 10% solution?
Answer by stanbon(75887)

how much water must be evaporated from 64oz of 2% solution to make a 10% solution?
----
Equation:
water - water = water
0.98*64 + x = 0.90(64+x)
------
Multiply thru by 100 and solve for "x":
98*64 + 100x = 90*64 + 90x
10x = -8*64
x = 51.2 oz (amt. that must be evaporated)
===============
Cheers,
Stan H.